OpenStudy (anonymous):

Given a vector function F = (3x)i + (4y)j -(7z)k, how can I find all vectors G such that F = curl(G)?

OpenStudy (anonymous):

curl(F)=0 curl(curl(G))=0 laplacian(G)-gard(Div(G))=0

OpenStudy (anonymous):

I know...but what's your point?

OpenStudy (anonymous):

actually i'm lost here...it seems to have many solutions..i tried to solve it using the help of Helmholt'z equation..

OpenStudy (anonymous):

yeah its a little difficult and bugging me. I'll keep working at it, thanks though

OpenStudy (anonymous):

\[\frac{d}{dy}G_z-\frac{d}{dz}G_y=3x\] and so on..you can also see that this type of problems generally have infinite solution..

OpenStudy (anonymous):

yeah i have those bunch of equations written down too! I'll keep at it.

OpenStudy (anonymous):

one thing i can suggest that take G(x,y,z)=X(x)Y(y)Z(z). this may help you

OpenStudy (anonymous):

yup ive done that and came up with those eqns. thanks tho.

OpenStudy (anonymous):

this problem seems to look like that you have given the mag. vector potential and you need to find the mag. field

OpenStudy (anonymous):

haha its exactly that if you apply the physics application.

OpenStudy (anonymous):

i gotta go...so bye..and sorry that i couldn't help you.

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

G(x,y,z) can not be equal to X(x)Y(y)Z(z) becuase G is a vector field

OpenStudy (anonymous):

he meant write it in the form of of G = Pi + Qj + Rk

OpenStudy (anonymous):

It can be G=4yzi-3xzj+3zk

OpenStudy (anonymous):

i need a general solution,

OpenStudy (anonymous):

G=4yzi-3xzj+Z(z)k also satisfying

OpenStudy (anonymous):

its easy to come up with a single example...i need for all vectors G

OpenStudy (anonymous):

OK wait..!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

it's a lengthy sol

OpenStudy (anonymous):

just describe in short what u did, ill understand

OpenStudy (anonymous):

Gx=X1(x)Y1(y)Z1(z) Gy=X2(x)Y2(y)Z2(z) Gz=X3(x)Y3(y)Z3(z)

OpenStudy (anonymous):

that means every component of G is a fcn of x,y,z... got that already :) proceed

OpenStudy (anonymous):

X3*Y3'*Z3-Z2'*X2*Y2=3x Z1' *Y1*X1-X3'*Y3*Z3=4y X2'*Y2*Z2-Y1'*X1*Z1=-7z

OpenStudy (anonymous):

now can you solve this system by your self??

OpenStudy (anonymous):

Here's what I have, cuz youre notation is a little strange.... Let G = P(x,y,z)i + Q(x,y,z)j + R(x,y,z)k So if curl(G) = F, 3x = (dR/dy)-(dQ/dz) 4y = (dP/dz)-(dR/dx) -7z = (dQ/dx)-(dP/dy)

OpenStudy (anonymous):

these are partial derivatives not exact. put Gx,Gy,Gz into that equation you wrote

OpenStudy (anonymous):

i know they are partials. ex dR/dy is my partial deriv of R with respect to y

OpenStudy (anonymous):

put Gx=X1(x)Y1(y)Z1(z) Gy=X2(x)Y2(y)Z2(z) Gz=X3(x)Y3(y)Z3(z) into 3x = (dR/dy)-(dQ/dz) 4y = (dP/dz)-(dR/dx) -7z = (dQ/dx)-(dP/dy)

OpenStudy (anonymous):

By Gx, do you mean the x-component of G? If so, then why do you write it as a product X1*Y1*Z1

OpenStudy (anonymous):

x component can be function of x,y,z, , you are dealing with vector fields!!

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

i dont understand your notation of X3*Y3*Z3 for example....

OpenStudy (anonymous):

you can write whatever you want I just wanted to write X1(x) is different to X3(x)

OpenStudy (anonymous):

ok fine i agree. but it doesn't necessarily have to be a product of xyz for Gx,Gy and Gz...

OpenStudy (anonymous):

for example??

OpenStudy (anonymous):

a sum

OpenStudy (anonymous):

(x1+x2)(y1+y2)=x1y1+x1y2+x2y1+x2y2

OpenStudy (anonymous):

your point?

OpenStudy (anonymous):

my point is nearly every expression can be factored so we can show in product, so in product nearly every possiblity is described

OpenStudy (anonymous):

hmm ok.

OpenStudy (anonymous):

even still, how would i solve that system you gave me before

OpenStudy (anonymous):

if you look at the first equation it's function of x so y3' z3=z2' y2=c so y3'/y2=z2'/z3=c so y3=c[integral]y2 dy+C1 and z2=c[integrall] z3 dz+C2 so first eqution becomes c(x3-x2)=3x I think you can proceed further by yourself :)

OpenStudy (anonymous):

hmm ill see. thanks tho

OpenStudy (anonymous):

I'm thinking about one more approach

OpenStudy (anonymous):

let G=(G1,G2,G3) G1=G1(x,y,z) so dG1=G1x dx+G1y dy+G1z dz from equations 4y = (dP/dz)-(dR/dx) -7z = (dQ/dx)-(dP/dy) we'll have G1=7zy+4yz+[integral] (partial derivativ w.r.t xG1)).dr similarly you can do for G2 and G3

OpenStudy (anonymous):

one more approach 3x = (dR/dy)-(dQ/dz) 4y = (dP/dz)-(dR/dx) -7z = (dQ/dx)-(dP/dy) let R=constant Q=-3xz+C(x,y) P=4yz+C(x,y).....